OpenStudy (anonymous):

How do I solve this a-b=6

OpenStudy (turingtest):

solve for what?

OpenStudy (anonymous):

Algebra

OpenStudy (turingtest):

no, I mean that you have two unknowns and one equation, so you can't get a number answer so do you want it solved for a or b ?

OpenStudy (anonymous):

Its my sons homework here is the problem the way its on the sheet. a-b=6 a=2b+7

OpenStudy (anonymous):

for both

OpenStudy (turingtest):

that changes things... you need as many equations as unknowns to solve a problem, so the other equation is essential a-b=6 a=2b+7 substitute the second expression for 'a' into the first equation: 2b+7-b=6 this can be solved for b once you have the value of b plug that number into the second equation, which gives you the value of a.

OpenStudy (anonymous):

so would i subtract the b or the 7 from the 6

OpenStudy (turingtest):

2b+7-b=6 collect like terms: b+7=6 then subtract 7 from both sides

OpenStudy (anonymous):

so b=(-1)

OpenStudy (turingtest):

yes, now plug that into b in the second equation a=2b+7 and you will get a

OpenStudy (anonymous):

is it a=5